\begin{thebibliography}{99}


 

\bibitem{said2}
 Albuquerque H., Barreiro E., Benayadi S., Boucetta M., Sánchez J.M.,
 Poisson algebras and symmetric Leibniz bialgebra structures on oscillator Lie algebras,
 Journal of Geometry and Physics, 160 (2021), 103939.

 

\bibitem{bai20}
 Bai C., Bai R., Guo L., Wu Y.,
Transposed Poisson algebras, Novikov-Poisson algebras, and 3-Lie algebras,  Journal of Algebra, 632 (2023), 535--566. 

\bibitem{angel}
Ballesteros A., Herranz F., 
Lie bialgebra quantizations of the oscillator algebra and their universal $R$-matrices, 
Journal of Physics A: Mathematical and Theoretical, 29 (1996), 15, 4307--4320.  

 
\bibitem{bfk23}
Beites P.,  Fern\'andez Ouaridi A., Kaygorodov I.,
The algebraic and geometric classification of  
transposed Poisson algebras, Revista de la Real Academia de Ciencias Exactas, Físicas y Naturales. Serie A. Matemáticas,  117 (2023), 2,    55.
 
 	\bibitem{bfk22}
		Beites P., 
		Ferreira B. L. M., 
		Kaygorodov I.,  
		Transposed Poisson structures,  arXiv:2207.00281.

	 

\bibitem{biggs}
Biggs R., Remsing C., 
Some remarks on the oscillator group, 
Differential Geometry and its Applications, 35 (2014), suppl., 199--209.


\bibitem{bm11}
Boucetta M.,  Medina A., 
Solutions of the Yang-Baxter equations on quadratic Lie groups: the case of oscillator groups, 
Journal of Geometry and Physics, 61 (2011), 12, 2309--2320.


\bibitem{cal}
Calvaruso G., Zaeim A., 
    On the symmetries of the Lorentzian oscillator group, 
    Collectanea Mathematica, 68 (2017), 1, 51--67.


 

\bibitem{crampe}
Crampé N., van de Vijver W.,  Vinet L., 
Racah problems for the oscillator algebra, the Lie algebra $\mathfrak{sl}_n$, and multivariate Krawtchouk polynomials, 
Annales Henri Poincaré, 21 (2020), 12, 3939--3971.
 

 
 \bibitem{dzhuma}
 Dzhumadil'daev  A.,
Weak Leibniz algebras and transposed Poisson algebras,
arXiv:2308.15018.  

 
 \bibitem{fer23}
 Fernández Ouaridi A.,
On the simple transposed Poisson algebras and Jordan superalgebras,
arXiv:2305.13848.   
 
 
 
\bibitem{FKL}
Ferreira B. L. M., Kaygorodov I., Lopatkin V., $\frac{1}{2}$-derivations of Lie algebras and transposed Poisson algebras, 
Revista de la Real Academia de Ciencias Exactas, Físicas y Naturales. Serie A. Matemáticas, 115 (2021), 3, 142. 
 
 
 	\bibitem{jac62}
Jacobson N.,  
Lie algebras, 
Interscience Tracts in Pure and Applied Mathematics, No. 10. Interscience Publishers (a division of John Wiley $\&$ Sons, Inc.), New York-London, 1962. ix+331 pp.

 
 	\bibitem{jawo}
		Jaworska-Pastuszak A., Pogorzały Z.,
		Poisson structures for canonical algebras,
		Journal of Geometry and Physics, 148 (2020), 103564.
		

\bibitem{kk21}
Kaygorodov I., Khrypchenko M., 
    Poisson structures on finitary incidence algebras, 
    Journal of  Algebra, 578 (2021), 402--420.
 
\bibitem{kk22}
Kaygorodov I., Khrypchenko M., 
		Transposed Poisson structures on Block Lie algebras and superalgebras, 
 Linear Algebra and Its Applications,
   656 (2023), 167--197.


\bibitem{kk23}
Kaygorodov I., Khrypchenko M., 
Transposed Poisson structures on Witt-type algebras,
 Linear Algebra and its Applications,    665  (2023),  196--210.	


\bibitem{kkg23}
Kaygorodov I., Khrypchenko M., 
Transposed Poisson structures on generalized Witt algebras and Block Lie algebras,    Results in Mathematics, 78 (2023),  5,  186.



\bibitem{KK7}
Kaygorodov I.,  Khrypchenko M.,
Transposed Poisson structures on the Lie algebra of upper triangular matrices, arXiv:2305.00727.


\bibitem{kkinc}
Kaygorodov I.,  Khrypchenko M.,
Transposed Poisson structures on   Lie incidence algebras, arXiv:2309.00332. 


\bibitem{klv22}
Kaygorodov I., Lopatkin V., Zhang Z.,
    Transposed Poisson structures on Galilean and solvable Lie algebras,  Journal of Geometry and Physics,
  187    (2023), 104781.


 

\bibitem{medina85}
Medina A.,  
Groupes de Lie munis de métriques bi-invariantes, 
Tohoku Mathematical Journal (2), 37 (1985), 4, 405--421.
  


   \bibitem{nw94}
Ndogmo J.,  Winternitz P., 
Solvable Lie algebras with abelian nilradicals, Journal of Physics A: Mathematical and Theoretical,   27 (1994),  2, 405--423.

   \bibitem{rw93}
Rubin J.,   Winternitz P., 
Solvable Lie algebras with Heisenberg ideals,
Journal of Physics A: Mathematical and Theoretical, 
 26 (1993), 5, 1123--1138.

   \bibitem{kms}
Sartayev B., 
Some generalizations of the variety of transposed Poisson algebras,
Communications in Mathematics, 32 (2024),  2, 54--61.
  
\bibitem{sw05} 
Šnobl L.,   Winternitz P., 
A class of solvable Lie algebras and their Casimir invariants,   
Journal of Physics A: Mathematical and Theoretical, 
38 (2005),  12, 2687--2700.



\bibitem{ytk} 
Yang Ya.,  Tang X.,  Khudoyberdiyev A.,	
Transposed Poisson structures on Schrodinger algebra in $(n+1)$-dimensional space-time, arXiv:2303.08180.

	\bibitem{YYZ07}
		Yao Y., Ye Y., Zhang P., 
		Quiver Poisson algebras, 
		Journal of  Algebra, 312 (2007), 2, 570--589.


    \bibitem{yh21}
Yuan L.,  Hua Q., 
$\frac{1}{2}$-(bi)derivations and transposed Poisson algebra structures on Lie algebras,
 Linear and Multilinear Algebra, 70 (2022),   22, 7672--7701.
 
  

\bibitem{fil1}  Zusmanovich P.,  
  On $\delta$-derivations of Lie algebras and superalgebras, Journal of Algebra, 324 (2010),  12, 3470--3486.
		
		
 
   
\end{thebibliography}

\documentclass[secthm,seceqn,amsthm,ussrhead,reqno, 12pt]{amsart}
\usepackage[utf8]{inputenc}
\usepackage[english]{babel}
\usepackage[symbol]{footmisc}
\usepackage{amssymb,amsmath,amsthm,amsfonts,xcolor,enumerate,hyperref,comment,longtable,cleveref}

\usepackage{times}
\usepackage{cite}
\usepackage{pdflscape}
\usepackage{ulem}
\usepackage[mathcal]{euscript}
\usepackage{tikz}
\usepackage{hyperref}
\usepackage{cancel}
\usepackage{stmaryrd}

\newtheorem{thrm}{Theorem} 

\newtheorem{corollary}[thrm]{Corollary}
\newtheorem{lem}[thrm]{Lemma}
\newtheorem{lemma}[thrm]{Lemma}

\newtheorem{proposition}[thrm]{Proposition}
\newtheorem{remark}[thrm]{Remark}

\newtheorem{defn}[thrm]{Definition}
\newtheorem{definition}[thrm]{Definition}

\crefrangeformat{equation}{#3(#1)#4--#5(#2)#6}

\crefname{thrm}{thrm}{Theorems}
\crefname{lem}{Lemma}{Lemmas}
\crefname{cor}{Corollary}{Corollaries}
\crefname{prop}{Proposition}{Propositions}
\crefname{proposition}{Proposition}{Propositions}
\crefname{defn}{Definition}{Definitions}
\crefname{exm}{Example}{Examples}
\crefname{rem}{Remark}{Remarks}
\crefname{section}{Section}{Sections}
\crefname{equation}{\unskip}{\unskip}
\crefname{enumi}{\unskip}{\unskip}

\DeclareMathOperator{\Ann}{Ann}

\newcommand{\gen}[1]{\langle #1\rangle}

\newcommand{\LL}{\mathfrak{L}}

\begin{document}

\noindent{\Large 
Transposed Poisson structures on 
 solvable and perfect Lie algebras}\footnote{
The  work is supported by 
FCT   UIDB/00212/2020, UIDP/00212/2020, 2022.02474.PTDC;
grant FZ-202009269, Ministry of higher education, science and innovations of the Republic of Uzbekistan.
} 

	\bigskip
	
	 \bigskip

\begin{center}	
	{\bf
		Ivan Kaygorodov\footnote{CMA-UBI, Universidade da Beira Interior, Covilh\~{a}, Portugal; 
  \    kaygorodov.ivan@gmail.com}	  \&
   Abror Khudoyberdiyev\footnote{ V.I.Romanovskiy Institute of Mathematics Academy of Science of Uzbekistan; National University of Uzbekistan; \ 
khabror@mail.ru}}   
\end{center}

	\bigskip
  
\noindent {\bf Abstract.}
{\it 
We described all transposed Poisson algebra structures on 
oscillator Lie algebras, i.e., on   one-dimensional 
solvable extensions of the $(2n+1)$-dimensional Heisenberg algebra; 
 on solvable Lie algebras with naturally graded filiform
nilpotent radical;
 on $(n+1)$-dimensional solvable   extensions of the $(2n+1)$-dimensional Heisenberg algebra; 
 and on $n$-dimensional solvable   extensions of the $n$-dimensional  algebra with the trivial multiplication.
We also gave an answer to one question on transposed Poisson algebras early posted   in 
a paper by Beites, Ferreira and Kaygorodov. 
Namely, we found a finite-dimensional Lie algebra with non-trivial $\frac{1}{2}$-derivations, 
but without non-trivial transposed Poisson algebra structures.

}

\bigskip

\noindent {\bf Keywords}: 
{\it Lie algebra, transposed Poisson algebra, 
$\delta$-derivation.}

\noindent {\bf MSC2020}: 17A30, 17B40, 17B61, 17B63. 

 \bigskip
  
\bigskip
\section*{Introduction}

	Since their origin in the 1970s in Poisson geometry, Poisson algebras have appeared in several areas of mathematics and physics, such as algebraic geometry, operads, quantization theory, quantum groups, and classical and quantum mechanics. One of the natural tasks in the theory of Poisson algebras is the description of all such algebras with fixed Lie or associative part~\cite{YYZ07,said2,jawo,kk21}.

	Recently, Bai, Bai, Guo and Wu have introduced a dual notion of the Poisson algebra~\cite{bai20}, called a \textit{transposed Poisson algebra}, by exchanging the roles of the two multiplications in the Leibniz rule defining a Poisson algebra. A transposed Poisson algebra defined this way not only shares some properties of a Poisson algebra, such as the closedness under tensor products and the Koszul self-duality as an operad but also admits a rich class of identities \cite{kms,bai20,fer23,bfk22}. It is important to note that a transposed Poisson algebra naturally arises from a Novikov-Poisson algebra by taking the commutator Lie algebra of its Novikov part \cite{bai20}.
 	Any unital transposed Poisson algebra is
	a particular case of a ``contact bracket'' algebra 
	and a quasi-Poisson algebra \cite{bfk22}.
 Each transposed Poisson algebra is a 
 commutative  Gelfand-Dorfman algebra \cite{kms}
 and it is also an algebra of Jordan brackets \cite{fer23}.
Transposed Poisson algebras are related to weak Leibniz algebras \cite{dzhuma}.
	In a recent paper by Ferreira, Kaygorodov, Lopatkin
	a relation between $\frac{1}{2}$-derivations of Lie algebras and 
	transposed Poisson algebras has been established \cite{FKL}. 	These ideas were used to describe all transposed Poisson structures 
	on  Witt and Virasoro algebras in  \cite{FKL};
	on   twisted Heisenberg-Virasoro,   Schr\"odinger-Virasoro  and  
	extended Schr\"odinger-Virasoro algebras in \cite{yh21};
	on Schr\"odinger algebra in $(n+1)$-dimensional space-time in \cite{ytk};
	on Witt type Lie algebras in \cite{kk23};
	on generalized Witt algebras in \cite{kkg23}; 
 Block Lie algebras in \cite{kk22,kkg23};
 on the Lie algebra of upper triangular matrices in \cite{KK7};
  and
on    Lie incidence algebras in \cite{kkinc}.
		Any complex finite-dimensional solvable Lie algebra was proved to admit a non-trivial transposed Poisson structure \cite{klv22}.
	The algebraic and geometric classification of $3$-dimensional transposed Poisson algebras was given in \cite{bfk23}.	
	For the list of actual open questions on transposed Poisson algebras, see \cite{bfk22}.	

 \medskip
 
The class of oscillator Lie algebras, which are the only noncommutative solvable Lie algebras that carry a bi-invariant Lorentzian metric \cite{medina85}, is a class of quadratic Lie algebras. The former algebras, which are the Lie algebras of oscillator Lie groups,
have aroused some interest \cite{biggs,said2,cal,crampe,bm11,angel}.
Ballesteros and  Herranz obtained all possible coboundary Lie bialgebras for the four-dimensional oscillator algebra \cite{angel}.
Biggs and  Remsing 
  examined the algebraic structure of the four-dimensional oscillator Lie algebra\cite{biggs}. Firstly, the adjoint orbits are determined (and graphed); these orbits turn out to be linearly isomorphic to the coadjoint orbits. This prompts the identification of a family of invariant scalar products (from which the Casimir functions can be recovered). They identified the group of automorphisms (resp. inner automorphisms); subsequently, they classified all linear subspaces. As a by-product, they arrived at classifications of the full-rank subspaces, the subalgebras, and the ideals.
 Based on the four-dimensional oscillator Lie algebra,
 Crampé, van de Vijver, and  Vinet introduced an associative algebra which they call the oscillator Racah algebra, and explain a relation between the representation theory of and multivariate Krawtchouk polynomials \cite{crampe}.
  In \cite{cal}, Calvaruso and Zaeim obtained a classification of all Ricci, curvature, Weyl, and matter collineations for the four-dimensional oscillator group equipped with a 1-parameter family of left-invariant Lorentzian metrics.
  Boucetta and Medina 
  determined the Lie bialgebra structures and the solutions of the classical Yang–Baxter equation on a generic class of oscillator Lie algebras \cite{bm11}.
   Albuquerque, Barreiro,  Benayadi,   Boucetta, and   Sánchez described all Poisson algebra structures and symmetric Leibniz bialgebra structures on generic oscillator Lie algebras \cite{said2}.
 The present paper is dedicated to studying transposed Poisson algebra structures on oscillator Lie algebras.
Namely, we have new examples of non-trivial ${\rm Hom}$-Lie algebra structures and transposed Poisson structures on oscillator Lie algebras.

\medskip
 
In Section \ref{tpaoscl}, we give the full description of all non-isomorphic transposed Poisson structures on generic oscillator Lie algebras, i.e., on   one-dimensional 
solvable extensions of the $(2n+1)$-dimensional Heisenberg algebra.
Section \ref{solv} is dedicated to a study of transposed Poisson algebra structures on some solvable algebras with a special type of nilpotent radical, obtained in some papers by Winternitz with co-authors \cite{sw05,nw94,rw93}.
Namely, we described all transposed Poisson algebra structures 
on   solvable Lie algebras with naturally graded filiform
nilpotent radical;
 on $(n+1)$-dimensional solvable   extensions of the $(2n+1)$-dimensional Heisenberg algebra; 
 and on $n$-dimensional solvable   extensions of the $n$-dimensional  algebra with the trivial multiplication.
In the last section, we also gave an answer to one question on transposed Poisson algebras early posted   in 
a paper by Beites, Ferreira, and Kaygorodov. 
Namely, we found a finite-dimensional Lie algebra with non-trivial $\frac{1}{2}$-derivations, 
but without non-trivial transposed Poisson algebra structures.

 \section{Preliminaries}\label{prem}
	All the algebras below will be over the complex field and all the linear maps will be $\mathbb C$-linear, unless otherwise stated. The notation $\gen{S}$ means the $\mathbb C$-subspace generated by $S$.

	\begin{defn}\label{tpa}
		Let ${\mathfrak L}$ be a vector space equipped with two nonzero bilinear operations $\cdot$ and $[-,-].$
		The triple $({\mathfrak L},\cdot,[-,-])$ is called a \textit{transposed Poisson algebra} if $({\mathfrak L},\cdot)$ is a commutative associative algebra and
		$({\mathfrak L},[-,-])$ is a Lie algebra that satisfies the following compatibility condition
		\begin{center}
			$2z\cdot [x,y]=[z\cdot x,y]+[x,z\cdot y].$
		\end{center}
	\end{defn}
	
	Transposed Poisson algebras were first introduced in a paper by Bai, Bai, Guo and Wu \cite{bai20}.
	
	\begin{defn}\label{tp-structures}
		Let $({\mathfrak L},[-,-])$ be a Lie algebra. A \textit{transposed Poisson algebra structure} on $({\mathfrak L},[-,-])$ is a commutative associative multiplication $\cdot$ on $\mathfrak L$ which makes $({\mathfrak L},\cdot,[-,-])$ a transposed Poisson algebra.
	\end{defn}

	\begin{defn}\label{12der}
		Let $({\mathfrak L}, [-,-])$ be an algebra with a multiplication $[-,-],$ $\varphi$ be a linear map
and $\phi$ be a bilinear map.
Then $\varphi$ is a $\frac{1}{2}$-derivation if it satisfies
\begin{center}
$\varphi[x,y]= \frac{1}{2} \big([\varphi(x),y]+ [x, \varphi(y)] \big);$
\end{center}
 $\phi$ is a $\frac{1}{2}$-biderivation if it satisfies
\begin{longtable}{crl}
$\phi([x,y],z)$&$=$&$ \frac{1}{2} \big( 
 [\phi(x,z),y] +[x,\phi(y,z)]\big),$\\ 
$\phi(x,[y,z])$&$=$&$ \frac{1}{2} \big([\phi(x,y),z]+ 
[y,\phi(x,z)] \big).$
\end{longtable}
 
	\end{defn}
	Observe that $\frac{1}{2}$-derivations are a particular case of $\delta$-derivations introduced by Filippov (see, for example,  \cite{fil1} and references therein).
 It is easy to see from Definition \ref{12der} that $[\LL,\LL]$ and $\Ann(\LL)$ are invariant under any $\frac 12$-derivation of $\LL$.

	\cref{tpa,12der} immediately implies the following key Lemma.
	\begin{lem}\label{glavlem}
		Let $({\mathfrak L},[-,-])$ be a Lie algebra and $\cdot$ a new binary (bilinear) operation on ${\mathfrak L}$. Then $({\mathfrak L},\cdot,[-,-])$ is a transposed Poisson algebra 
		if and only if $\cdot$ is commutative and associative and for every $z\in{\mathfrak L}$ the multiplication by $z$ in $({\mathfrak L},\cdot)$ is a $\frac{1}{2}$-derivation of $({\mathfrak L}, [-,-]).$
	\end{lem}
	
	The basic example of a $\frac{1}{2}$-derivation is the multiplication by a field element.
	Such $\frac{1}{2}$-derivations will be called \textit{trivial}. 
	
	\begin{thrm}\label{princth}
		Let ${\mathfrak L}$ be a Lie algebra without non-trivial $\frac{1}{2}$-derivations.
		Then all transposed Poisson algebra structures on ${\mathfrak L}$ are trivial.
	\end{thrm}

	 Let $\cdot$ be a transposed Poisson algebra structure on a Lie algebra $({\mathfrak L}, [-,-])$. 
	 Then any automorphism $\phi$ of $({\mathfrak L}, [-,-])$ induces the transposed Poisson algebra structure $*$ on $({\mathfrak L}, [-,-])$ given by
	 \begin{align*}
		 x*y=\phi\big(\phi^{-1}(x)\cdot\phi^{-1}(y)\big),\ \ x,y\in{\mathfrak L}.
		 \end{align*}
	 Clearly, $\phi$ is an isomorphism of transposed Poisson algebras $({\mathfrak L},\cdot,[-,-])$ and $({\mathfrak L},*,[-,-])$.

\section{Transposed Poisson algebra structures on oscillator Lie algebras }\label{tpaoscl}

\begin{definition}
The oscillator Lie algebra $\mathfrak{L_{\lambda}}$ is a $(2n+2)$-dimensional Lie algebra with its canonical basis $\mathbb{B} = \{ e_{-1}, e_{0}, e_j, \check{e}_j \}_{j=1,\ldots,n}$  
and the Lie brackets are given by
 $$[e_{-1},e_j] = \lambda_j\check{e}_j, \ \ \  [e_{-1},\check{e}_j] = -\lambda_je_j, \ \ \ [e_j,\check{e}_j] = e_0,$$
for $j = 1, \ldots, n$ and $\lambda = (\lambda_1, \ldots, \lambda_n) \in \mathbb{R}^{n}$ with $0 < \lambda_1 \leq \ldots \leq \lambda_n$.
\end{definition}

\begin{proposition}\label{dmatrix} 
A linear map $\varphi : \mathfrak{L}_{\lambda} \rightarrow \mathfrak{L}_{\lambda}$ is a $\frac{1}{2}$-derivation of the algebra $(\mathfrak{L}_{\lambda}, [-,-])$ if and only if 
 
\begin{longtable}{lcl}
$\varphi(e_{-1})$ &$ =$&$ \gamma e_{-1} + \mu e_0 - \sum\limits_{j=1}^{n} 2\lambda_j \alpha_{j} e_j - \sum\limits_{j=1}^{n} 2 \lambda_j \beta_{j} \check{e}_j$,\\
$\varphi(e_0)$ &$ =$& $ \gamma e_0$,\\
$\varphi(e_j)$ & $=$ & $\alpha_{j} e_0 + \gamma e_j, \ \ \ (j=1, \ldots, n)$,\\
$\varphi(\check{e}_j)$ & $ =$ &$ \beta_{j} e_0 + \gamma \check{e}_{j}, \ \ \ (j=1, \ldots, n).$
\end{longtable}
\end{proposition}
\begin{proof}
For each $\ell \in \mathbb{B},$  we consider 
\begin{center}
$\varphi(\ell) = \varphi(\ell)_{-1}e_{-1} + \varphi(\ell)_{0}e_0 + \sum\limits_{k=1}^{n} (\varphi(\ell)_{k}e_k +  \varphi(\ell)_{\check{k}}\check{e}_k)$.
\end{center}

Since $[\LL,\LL]$ and $\Ann(\LL)$ are invariant under any $\frac 12$-derivation, we obtain that $\varphi(e_0) = \gamma e_0$ and $\varphi(e_j)_{-1} = \varphi(\check{e}_j)_{-1} =0.$

Now from $e_0 = [e_j,\check{e}_j]$ for each $j \in \{1, \ldots, n\},$ we get
$$\varphi(e_0) = \varphi([e_j,\check{e}_j]) = \frac{1}{2}([\varphi(e_j),\check{e}_j] + [e_j,\varphi(\check{e}_j)]) =
\frac{1}{2}(\varphi(e_j)_{j} + \varphi(\check{e}_j)_{\check{j}}e_0),$$
which implies  $\varphi(e_j)_j+\varphi(\check{e}_j)_{\check{j}} = 2\gamma.$

Since $[e_{-1},e_j] = \lambda_j \check{e}_j$ and $[e_{-1},\check{e}_j] = -\lambda_je_j$ then
\begin{longtable}{rcl}
$\varphi(\lambda_j\check{e}_j)$ &$ =$&$ \varphi([e_{-1},e_j]) = \frac{1}{2}([\varphi(e_{-1}),e_j] + [e_{-1},\varphi(e_j)])$ \\
&$=$& $\frac{1}{2} \lambda_j \varphi(e_{-1})_{-1} \check{e}_j - \frac{1}{2} \varphi(e_{-1})_{\check{j}} e_0   +\frac{1}{2} \sum\limits_{k=1}^n\Big(\lambda_k \varphi(e_j)_{k} \check{e}_k -\lambda_k \varphi(e_j)_{\check{k}} e_k\Big),$
\end{longtable}
and
\begin{longtable}{rcl}
$-\varphi(\lambda_je_j)$&$ =$&$ \varphi([e_{-1},\check{e}_j]) = \frac{1}{2}([\varphi(e_{-1}),\check{e}_j] + [e_{-1},\varphi(\check{e}_j)])$ \\
&$=$&$ -\frac{1}{2} \lambda_j \varphi(e_{-1})_{-1} e_j + \frac{1}{2} \varphi(e_{-1})_{j} e_0 +\frac{1}{2} \Big(\sum\limits_{k=1}^n \lambda_k \varphi(\check{e}_j)_{k} \check{e}_k -\lambda_k \varphi(\check{e}_j)_{\check{k}} e_k\Big).$
\end{longtable}

Hence,  for any $j \in \{1, \ldots, n\},$ we have 
$$\varphi(e_{-1})_{-1} = \varphi(e_j)_j= \varphi(\check{e}_j)_{\check{j}} =\gamma, \quad \varphi(e_{-1})_j=-2\lambda_j \varphi(e_j)_0, \quad  \varphi(e_{-1})_{\check{j}}=-2\lambda_j \varphi(e_{\check{j}})_0,$$ 
and
\begin{equation}\label{eq1}
 2\lambda_j \varphi(\check{e}_j)_k   = - \lambda_k \varphi(e_j)_{\check{k}}, \quad 2\lambda_j  \varphi(e_j)_{\check{k}}  = - \lambda_k \varphi(\check{e}_j)_k, \quad 1 \leq j, k \leq n,
\end{equation}

\begin{equation}\label{eq2}
 2\lambda_j \varphi(\check{e}_j)_{\check{k}}   =  \lambda_k \varphi(e_j)_k, \quad 2\lambda_j  \varphi(e_j)_k  =  \lambda_k \varphi(\check{e}_j)_{\check{k}}, \quad 1 \leq j, k (j \neq k) \leq n.
\end{equation}

Finally, from
\begin{longtable}{ccccc}
$0$& $=$ & $\varphi([e_j, e_k])$ &$=$ & $\frac{1}{2}([\varphi(e_j),e_k] + [e_j,\varphi(e_k)]),$\\
$0$ &$=$ & $\varphi([e_j, \check{e}_k])$ &$=$ & $\frac{1}{2}([\varphi(e_j),\check{e}_k] + [e_j,\varphi(\check{e}_k)]),$\\
$0$ &$=$ & $\varphi([\check{e}_j, \check{e}_k])$ &$=$ & $\frac{1}{2}([\varphi(\check{e}_j),\check{e}_k] + [\check{e}_j,\varphi(\check{e}_k)]),$
\end{longtable}
for $1 \leq j, k (j \neq k) \leq n,$ we derive that 
\begin{equation}\label{eq3}
 \varphi(e_j)_{\check{k}}   = \varphi(e_k)_{\check{j}}, \quad 
 \varphi(e_j)_k   = - \varphi(\check{e}_k)_{\check{j}}, \quad 
 \varphi(\check{e}_j)_k   = \varphi(\check{e}_k)_j, 
\end{equation}

It is not difficult to deduce from \eqref{eq1}, \eqref{eq2} and \eqref{eq3} that 
 \begin{center} $ \varphi(e_j)_t=0$ for $t\in \{1,  \ldots, n, \check{1}, \ldots, \check{n}\} \setminus \{ j\},$\end{center}
\begin{center} $ \varphi(e_{\check{j}})_t=0$ for $t \in \{1,  \ldots, n, \check{1}, \ldots, \check{n}\} \setminus \{ \check{j}\}.$ \end{center}


Hence, our map $\varphi$ has the following type:
\begin{longtable}{rcl}
$\varphi(e_{-1})$ &$ =$&$ \gamma e_{-1} + \mu e_0 - \sum\limits_{j=1}^{n} 2\lambda_j \alpha_{j} e_j - \sum\limits_{j=1}^{n} 2 \lambda_j \beta_{j} \check{e}_j$,\\
$\varphi(e_0)$ &$ =$& $ \gamma e_0$,\\
$\varphi(e_j)$ & $=$ & $\alpha_{j} e_0 + \gamma e_j, \ \ \ (j=1, \ldots, n)$,\\
$\varphi(\check{e}_j)$ & $ =$ &$ \beta_{j} e_0 + \gamma \check{e}_{j}, \ \ \ (j=1, \ldots, n).$
\end{longtable}
It is easy to see that all linear maps given by the present way are $\frac{1}{2}$-derivations.
It finishes the proof.

\end{proof}

 Filippov proved that each $\delta$-derivation ($\delta\neq0,1$) gives a non-trivial ${\rm Hom}$-Lie algebra structure \cite[Theorem 1]{fil1}.
 Hence, by Proposition \ref{dmatrix}, we have the following corollary.

  \begin{corollary}
  Each oscillator Lie algebra admits a non-trivial ${\rm Hom}$-Lie algebra structure.
 \end{corollary}

\begin{thrm}\label{theodot}
Let $(\mathfrak{L}_{\lambda},  [-,-])$ be the oscillator Lie algebra. Then every symmetric $\frac{1}{2}$-biderivation gives a transposed Poisson structure.
If $(\mathfrak{L}_{\lambda}, \cdot, [-,-])$ is a transposed  Poisson algebra, then  $(\mathfrak{L}_{\lambda}, \cdot)$ has  the following  multiplication:
\begin{longtable}{lcl}
$e_{-1} \cdot e_{-1}$ & $ =$ &$ \gamma  e_{-1}+ \mu e_0 - 2\sum\limits_{k=1}^{n} \lambda_k (\alpha_{k} e_k +\beta_{k} \check{e}_k),$\\
$ e_{-1} \cdot e_{0}$&$ =$&$ \gamma e_0,$\\
$e_{-1} \cdot e_{j}$ & $ =$ & $ \alpha_{j} e_0 + \gamma e_{j}, $\\
$e_{-1} \cdot \check{e}_{j}$ & $ =$ & $ \beta_{j} e_0 + \gamma \check{e}_{j},$\\
$e_j \cdot e_j$ & $ =$ & $ \check{e}_j \cdot \check{e}_j = -\frac{\gamma}{2\lambda_j} e_0  .$
\end{longtable}
$(\mathfrak{L}_{\lambda}, \cdot, [-,-])$ is a Poisson algebra if and only if
$(\gamma, \alpha_1, \ldots \alpha_n, \beta_1, \ldots, \beta_n)=0.$
\end{thrm}

\begin{proof}
We aim to describe the multiplication $\cdot.$
By Lemma \ref{glavlem}, 
for each  $\ell \in \mathbb{B}$, there is a related 
$\frac{1}{2}$-derivation $\varphi_{\ell}$ of $(\mathfrak{L}_{\lambda}, [-,-])$
such that $\varphi_{\ell_1}(\ell_2 )=\ell_1 \cdot \ell_2 = \varphi_{\ell_2}(\ell_1),$ 
with $\ell_1, \ell_2 \in \mathbb{B}.$
By Proposition \ref{dmatrix}, for each  $\ell \in \mathbb{B}$, we have

\begin{longtable}{rcl}
$\varphi_{\ell}(e_{-1})$ & $ =$ &$\gamma^{\ell} e_{-1} + \mu^{\ell} e_0 - \sum\limits_{k=1}^{n} 2\lambda_k \alpha^{\ell}_{k} e_k - \sum\limits_{k=1}^{n} 2 \lambda_k \beta^{\ell}_{k} \check{e}_k$,\\
$\varphi_{\ell}(e_0)$ & $ =$& $  \gamma^{\ell} e_0$,\\
$\varphi_{\ell}(e_j) $&$=$&$ \alpha^{\ell}_{j} e_0 + \gamma^{\ell} e_j, \ \ \ (j=1, \ldots, n)$,\\
$\varphi_{\ell}(\check{e}_j)$ & $ =$&$ \beta^{\ell}_{j} e_0 + \gamma^{\ell} \check{e}_{j}, \ \ \ (j=1, \ldots, n)$.
\end{longtable}
In the first, for all $i,j \in \{1, \ldots, n\},$ we have 
\begin{longtable}{rcccccl}

$\alpha^{e_{i}}_{j} e_0 + \gamma^{e_{i}} e_j $&$=$&$ \varphi_{e_{i}}(e_j) $&$=$&$ \varphi_{e_{j}}(e_i) $&$=$&$ \alpha^{e_{j}}_{i} e_0 + \gamma^{e_{j}} e_i$,\\
$\alpha^{\check{e}_i}_{j} e_0 + \gamma^{\check{e}_i} e_j$&$ =$&$ \varphi_{\check{e}_i}(e_j) $&$=$&$  \varphi_{e_{j}}(\check{e}_i)$&$ =$&$ \beta^{e_{j}}_{i} e_0 + \gamma^{e_{j}} \check{e}_{i}$,\\
$\beta^{\check{e}_i}_{j} e_0 + \gamma^{\check{e}_i} \check{e}_{j}$&$ =$&$ \varphi_{\check{e}_i}(\check{e}_j)$&$ = $&$\varphi_{\check{e}_j}(\check{e}_i) $&$ =$&$\beta^{\check{e}_j}_{i} e_0 + \gamma^{\check{e}_j} \check{e}_{i}.$
\end{longtable}

Hence 
$\beta^{\check{e}_j}_{i} = \beta^{\check{e}_i}_{j}$, 
$\alpha^{e_{i}}_{j}=\alpha^{e_{j}}_{i} $, 
$\alpha^{\check{e}_i}_{j} = \beta^{e_{j}}_{i}$ and 
$\gamma^{e_{j}} = \gamma^{\check{e}_j} =  0$.

Second, we have
$$
\gamma^{e_{-1}}  e_0 = \varphi_{e_{-1}}(e_0) =\varphi_{e_{0}}(e_{-1}) =  \gamma^{e_{0}} e_{-1}+\mu^{e_0}e_0
-2 \sum\limits_{k=1}^n \lambda_k (\alpha_k^{e_0} e_k+\beta_k^{e_0} \check{e}_k),$$
which implies $ \alpha^{e_0}_j=\beta^{e_0}_j= \gamma^{e_{0}} = 0$ and $\gamma^{e_{-1}}=\mu^{e_0}$.

Also,  for all $ j \in \{1, \ldots, n\},$ we have 
\begin{longtable}{rcccccl}
$\alpha^{e_{-1}}_{j} e_0 + \gamma^{e_{-1}} e_j$&$ = $&$ \varphi_{e_{-1}}(e_j) $&$ = $&$ \varphi_{e_{j}}(e_{-1}) $&$ = $&$    
\mu^{e_{j}} e_0 - 2\sum\limits_{k=1}^{n} \lambda_k (\alpha^{e_{j}}_{k} e_k+\beta^{e_{j}}_{k} \check{e}_k),$\\
$\beta^{e_{-1}}_{j} e_0 + \gamma^{e_{-1}} \check{e}_{j} $&$ = $&$ \varphi_{e_{-1}}(\check{e}_j)$&$ = $&$ \varphi_{\check{e}_j}(e_{-1}) $&$ = $&$  
 \mu^{\check{e}_j} e_0 - 2\sum\limits_{k=1}^{n}  \lambda_k( \alpha^{\check{e}_j}_{k} e_k +  \beta^{\check{e}_j}_{k} \check{e}_k),$
\end{longtable}
which implies
\begin{longtable}{llll}
$\gamma^{e_{-1}} =  -2\lambda_j \alpha^{e_j}_j,$ & 
$ \alpha^{e_j}_j=\beta^{\check{e}_j}_j ,$ &  
$\alpha^{\check{e}_j}_{k} = 0,$ & 
$\beta^{e_j}_k = 0,$ \\ 

$\alpha^{e_j}_k = 0 \ (j\neq k),$ & 
$\beta^{\check{e}_j}_{k} = 0  \ (j\neq k),$ & 
$\alpha^{e_{-1}}_j = \mu^{e_j},$ & 
$\beta^{e_{-1}}_j = \mu^{\check{e}_j}$.
\end{longtable}

It follows that, for all $j \in \{1, \ldots, n\}$,   the multiplication $\cdot$ can be written as
\begin{longtable}{lcl}
$e_{-1} \cdot e_{-1}$ & $ =$ &$ \gamma  e_{-1}+ \mu e_0 - 2\sum\limits_{k=1}^{n} \lambda_k (\alpha_{k} e_k +\beta_{k} \check{e}_k),$\\
$ e_{-1} \cdot e_{0}$&$ =$&$ \gamma e_0,$\\
$e_{-1} \cdot e_{j}$ & $ =$ & $ \alpha_{j} e_0 + \gamma e_{j}, $\\
$e_{-1} \cdot \check{e}_{j}$ & $ =$ & $ \beta_{j} e_0 + \gamma \check{e}_{j},$\\
$e_j \cdot e_j$ & $ =$ & $ \check{e}_j \cdot \check{e}_j = -\frac{\gamma}{2\lambda_j} e_0  .$
\end{longtable}


It is easy to see that $\cdot$ is a commutative associative multiplication.

\end{proof}

\subsection*{Isomorphism problem for transposed Poisson algebras on generic oscillator Lie algebras}
An oscillator Lie algebra is called generic if 
$0< \lambda_1 < \ldots  < \lambda_n$ and $\lambda_i+\lambda_j \neq \lambda_k$ for all $1\leq i < j<k \leq n.$ 
It is easy to see that
$(\mathfrak{L}_{ \{ \lambda_1, \lambda_2, \ldots, \lambda_n\}}, [-,-])$ is isomorphic to
$(\mathfrak{L}_{ \left\{ 1, \frac{\lambda_2}{\lambda_1}, \ldots, \frac{\lambda_n}{\lambda_1}\right\}}, [-,-]).$ 
Hence, we will consider only the case $\lambda_1=1.$

 To prove the main result of this part of the paper, we should give the following useful Lemma.
 
 \begin{lemma}\label{autgen}
 Let $\phi$ be an   automorphism of $(\mathfrak{L}_{\lambda}, [-,-])$.
 Then $\phi$ has the following types:
 \begin{longtable}{rcl}
$ \phi(e_{-1})$&$=$&$  \pm e_{-1} + \nu e_0+\sum\limits_{i=1}^n \nu_i e_i+\sum\limits_{i=1}^n \check{\nu}_i \check{e}_i,$\\ 
$ \phi(e_0)$&$=$&$\pm \xi e_0, $\\ 
$ \phi(e_i)$&$=$&$  \frac{ \check{\nu}_i\check{\mu}_i \mp \nu_i \mu_i}{\lambda_i}e_0+  \mu_i e_i \mp \check{\mu}_i \check{e}_i,$\\ 
$ \phi(\check{e}_i)$&$=$&$   \frac{ -\check{\nu}_i\check{\mu}_i \mp \nu_i \mu_i}{\lambda_i}e_0+  \check{\mu}_i e_i \pm \mu_i \check{e}_i,$\\ 
      \end{longtable}
  where $\xi= \mu_i^2 +\check{\mu}_i^2.$  
 \end{lemma}
 \begin{proof}
 Let us say that 
 \begin{center}
     $\phi(e_m)= \alpha^{m}_{-1} e_{-1}+\alpha^{m}_{0} e_{0}+\sum\limits_{i=1}^{n}(\alpha_i^m e_i+\beta_i^m \check{e}_i),$ where $m\in \{-1,0,1, \ldots, n\}.$
      \end{center}

Since $\Ann(\LL)$ and $[\LL,\LL]$ are invariant under any automorphism, we have that 
$$\alpha_{-1}^{0} =\alpha_{i}^{0}=\beta_{i}^{0}=0, \quad \alpha_{i}^{-1}=0, \quad 1 \leq i \leq n.$$

Then, for each $j \in \{1,\ldots, n\},$ we have
\[\lambda_j^2 \phi(e_j)=-[\phi(e_{-1}),[\phi(e_{-1}),\phi(e_{j})]]=
(\alpha^{-1}_{-1})^2\sum\limits_{i=1}^n\lambda_i^2(\alpha_i^j e_i+\beta_i^j \check{e}_j)
-\alpha^{-1}_{-1}\sum\limits_{i=1}^n(\alpha_{i}^{-1}\alpha_i^j+\beta_i^{-1}\beta_i^j)e_0.
\]

Since $\lambda_i\neq \lambda_j,$ we have that 
$$\alpha_{-1}^{-1}=\pm1,  \quad \alpha^j_0 = -\frac{\alpha_{-1}^{-1}}{\lambda_j}(\alpha_{j}^{-1}\alpha_j^j+\beta_j^{-1}\beta_j^j), \quad \alpha_i^j = \beta^j_i =0, \quad 1 \leq i (i\neq j) \leq n.$$

Thus, we obtain $$\phi(e_j)=-\frac{\alpha_{-1}^{-1}}{\lambda_j}(\alpha_{j}^{-1}\alpha_j^j+\beta_j^{-1}\beta_j^j)e_0+\alpha_j^je_j+\beta^j_j\check{e}_j$$

In a similar way, considering 
$\lambda_j^2 \phi(\check{e}_j)=-[\phi(e_{-1}),[\phi(e_{-1}),\phi(\check{e}_{j})]],$ we have 
$$\phi(\check{e}_j)=-\frac{\alpha_{-1}^{-1}}{\lambda_j}(\alpha_{j}^{-1}\mu_j^j+\beta_j^{-1}\nu_j^j)e_0+\mu_j^je_j+\nu^j_j\check{e}_j.$$

Finally, using the equalities $$\phi[e_{i},\check{e}_{i}] = [\phi(e_{i}),\phi(\check{e}_{i})],\quad \phi[e_{-1},{e}_{i}] = [\phi(e_{-1}),\phi({e}_{i})], \quad \phi[e_{i},\check{e}_{i}] = [\phi(e_{i}),\phi(\check{e}_{i})],$$
we derive that $\phi$ has the form from our statement.
\end{proof}

For our main theorem from this part, we need to define the following set
\begin{center}
    $\mathbb C_{>0}= \{ a+ b {\rm  i}  \}_{a>0} \cup \{  b{\rm  i} \}_{b\geq0}.$ 
\end{center}

\begin{thrm}
Let $(\mathfrak{L}_{\lambda},  [-,-])$ be the generic oscillator Lie algebra. If $(\mathfrak{L}_{\lambda}, \cdot, [-,-])$ is a transposed  Poisson algebra structure on $(\mathfrak{L}_{\lambda}, [-,-])$, 
then  $(\mathfrak{L}_{\lambda}, \cdot)$ 
has the following multiplication.
\begin{enumerate}[(A)]
    \item If  $(\mathfrak{L}_{\lambda}, \cdot)$ is non-nilpotent, then  $\gamma\neq0$ and we have:
\begin{center}
$e_{-1} \cdot e_{-1}=\gamma  e_{-1}, \
e_{-1} \cdot e_{0}= \gamma e_0, \
e_{-1} \cdot e_{j}= \gamma e_{j}, \
e_{-1} \cdot \check{e}_{j}=  \gamma \check{e}_{j}, \
e_j \cdot e_j= \check{e}_j \cdot \check{e}_j = -\frac{\gamma}{2\lambda_j} e_0,$
\end{center}
corresponding transposed Poisson algebra  $(\mathfrak{L}_{ \{\lambda, \gamma\}}, \cdot, [-,-])$ is  isomorphic only to  $(\mathfrak{L}_{ \{\lambda, -\gamma\}}, \cdot, [-,-])$; 
all these algebras are non-Poisson.

  \item If  $(\mathfrak{L}_{\lambda}, \cdot)$ is  nilpotent, then
  there is a sequence $\{ \beta_q \in \mathbb C_{>0} \}_{ 1\leq q \leq n},$ 
  where the first nonzero element is equal to $1,$
  and $(\mathfrak{L}_{\lambda}, \cdot)$ has the  following multiplication table:
\begin{enumerate}
    \item 
$e_{-1} \cdot e_{-1}=   e_{0} -2\sum\limits_{j=1}^n \lambda_j \beta_j \check{e}_{j}, \
e_{-1} \cdot \check{e}_{j}=  \beta_j \check{e}_{j},$

    \item 
$e_{-1} \cdot e_{-1}=    -2\sum\limits_{j=1}^n \lambda_j \beta_j \check{e}_{j}, \
e_{-1} \cdot \check{e}_{j}=  \beta_j \check{e}_{j}.$
\end{enumerate}
The corresponding transposed Poisson algebras  
$(\mathfrak{L}_{\{\lambda, \beta\}}, \cdot, [-,-])$ are not isomorphic;
and they are non-Poisson if there exists $\beta_j \neq0$.
\end{enumerate}

\end{thrm}

\begin{proof}
Let us consider the transposed Poisson structure  $(\mathfrak{L}_{\lambda}, \cdot , [-,-])$ 
on the generic oscillator Lie $(\mathfrak{L}_{\lambda},  [-,-])$ 
and $\cdot$ is given as in Theorem \ref{theodot}.
Under the action of a positive automorphism of the Lie algebra $(\mathfrak{L}_{\lambda},  [-,-]),$
given in Lemma \ref{autgen}, we rewrite the multiplication table of $(\mathfrak{L}_{\lambda}, \cdot)$ by the following way:

\begin{longtable}{lcl}
$e_{-1} \cdot e_{-1} $&$=$&$ \gamma e_{-1}+\left( \frac{\mu+\gamma \nu}{\xi}+\sum\limits_{i=1}^{n}\frac{\gamma(\nu_i^2+\check{\nu}_i^2)}{2 \lambda_i \xi}\right)e_0+$\\
& & \multicolumn{1}{r}{$\sum\limits_{k=1}^n\left( \frac{\gamma( \mu_k\nu_k- \check{\mu}_k\check{\nu}_k)-2\lambda_k(\alpha_k  \mu_k-\beta_k  \check{\mu}_k) }{\xi}\right)e_k+
\sum\limits_{k=1}^n\left( \frac{\gamma( \check{\mu}_k\nu_k+ \mu_k\check{\nu}_k)-2\lambda_k(\alpha_k  \check{\mu}_k+\beta_k  \mu_k) }{\xi}\right)\check{e}_k$}\\
$e_{-1} \cdot e_{0} $&$=$&$\gamma e_0$\\

$e_{-1} \cdot e_{k} $&$=$&$\left( \frac{\gamma( \mu_k\nu_k- \check{\mu}_k\check{\nu}_k)-2\lambda_k(\alpha_k  \mu_k-\beta_k  \check{\mu}_k) }{2\lambda_k\xi}\right)e_0 +\gamma e_k$\\

$e_{-1} \cdot \check{e}_{k} $&$=$&$\left( \frac{\gamma( \check{\mu}_k\nu_k+ \mu_k\check{\nu}_k)-2\lambda_k(\alpha_k  \check{\mu}_k+\beta_k  \mu_k) }{2\lambda_k\xi}\right)e_0 +\gamma \check{e}_k$\\
$e_k \cdot e_k$ & $ =$ & $ \check{e}_k \cdot \check{e}_k = -\frac{\gamma}{2\lambda_k} e_0  .$
\end{longtable}
Hence, we will consider two different situations.
\begin{enumerate} 
    \item[$\gamma\neq 0.$]
    By taking 
\begin{center}    $\nu_k=\frac{2\alpha_k \lambda_k}{\gamma}, $
    $\check{\nu}_k=\frac{2\beta_k \lambda_k}{\gamma} $ and
    $\nu= -\sum\limits_{i=1}^{n}\frac{\gamma(\nu_i^2+\check{\nu}_i^2)}{2 \gamma \lambda_i  }-\frac{\mu}{\gamma},$
\end{center}
we have a multiplication table from part $(A)$ in our statement.
It is easy to see that, under the action of a positive automorphism of the Lie algebra $(\mathfrak{L}_{\lambda},  [-,-]),$
given in Lemma \ref{autgen}, all  multiplications $\cdot$ give  non-isomorphic transposed Poisson structures. 
Under the action of a negative automorphism given in Lemma \ref{autgen}, we have the isomorphism between 
structures corresponding to $\gamma$ and $-\gamma.$
Insofar, $(\mathfrak{L}_{\lambda},   \cdot )$ is non-nilpotent if and only if $\gamma \neq 0$.

  \item[$\gamma= 0.$] In this case,     $(\mathfrak{L}_{\lambda},   \cdot )$ is a nilpotent algebra. Note that, 
  \begin{enumerate}[$\bullet$]
      \item 
        If   $\{ \alpha_k \}_{1 \leq k \leq n}=0,$ then transposed Poisson structures corresponding to 
        $\{\beta_1, \ldots, \beta_j, \ldots, \beta_n \}$ and 
        $\{\beta_1, \ldots, -\beta_j, \ldots, \beta_n \}$ are isomorphic.

      \item 
  If $(\alpha_k,\beta_k)\neq (0, 0)$ then after an action of a suitable automorphism, we can suppose that $\alpha_k\neq 0$ and $\beta_k\neq0.$ 
  In the case  $(\alpha_k,\beta_k)=(0, 0)$ after each automorphism action, we will take the same situation  $(\alpha_k,\beta_k)=(0, 0).$
   Let $t$ be the maximal number such that $(\alpha_1,\beta_1)=\ldots=(\alpha_{t-1},\beta_{t-1})=(0, 0).$
  \end{enumerate}
  We should consider two subcases.
  \begin{enumerate}

    \item [$\mu\neq0.$] 
    By taking 
\begin{center}
    $\xi=\mu ,$ $\check{\mu}_t=\alpha_t$ and $\mu_k=\beta_k$ for all $k $ such that $\alpha_k\beta_k\neq0,$ 
\end{center}
we obtain that $(\mathfrak{L}_{\lambda}, \cdot)$ has a multiplication table from the part $(B.a)$ of our statement.

   \item [$\mu=0.$] 
    By taking 
\begin{center}
     $\check{\mu}_t=\alpha_t$ and $\mu_k=\beta_k$ for all $k $ such that $\alpha_k \beta_k\neq0,$ 
\end{center}
we obtain that $(\mathfrak{L}_{\lambda}, \cdot)$ has a multiplication table from the part $(B.b)$ of our statement.

  \end{enumerate}
\end{enumerate}
\end{proof}

\section{Transposed Poisson structures on  finite-dimensional solvable Lie algebras}\label{solv}

It is proven that any finite-dimensional solvable Lie algebra over an algebraically closed field of zero characteristic admits non-trivial $\frac12$-derivations \cite{klv22}. For the solvable Lie algebra $\mathfrak L$ with non-zero annihilator, it is constructed non-trivial $\frac12$-derivation as $\varphi(x) = [x, \omega_0],$
by the element of $\omega_0 \in \operatorname{Ann}_{[\mathfrak{L},\mathfrak{L}]}([\mathfrak{L},\mathfrak{L}])$ such that $[x,\omega_0] =\lambda_x\omega_0$ for all $x \in \mathfrak{L},$ where $\lambda_x \neq 0$ for some $x.$ Thus, the dimension of the $\frac12$-derivations space of any solvable Lie algebra at least equal to two. 
In the following example, we give a solvable Lie algebra in which the dimension of the $\frac12$-derivations space is equal to two.

\subsection{Transposed Poisson structures on solvable Lie algebras with naturally graded filiform nilpotent radical}
All solvable Lie algebras with a naturally graded filiform nilpotent radical were found in \cite{sw05}.
Now we consider the  solvable Lie algebra $\mathfrak{s}_{n,2}$ with the following multiplication table:
\begin{longtable}{lcll lcll}
$[e_i, e_1] $&$=$&$ e_{i+1},$ \ &   $2 \leq i \leq n-1;$ &
$[e_1, x_1] $&$=$&$ e_1,$ \\
$[e_i, x_1] $&$=$&$ (i-2)e_i,$\  & $3 \leq i \leq n;$ &
$[e_i, x_2] $&$=$&$ e_i,$ \ &  $2 \leq i \leq n.$
\end{longtable}

Note that $\mathfrak{s}_{n,2}$ is a  solvable Lie algebra with nilpotent radical is the naturally graded filiform Lie algebra
$$\mathfrak{n}_{n,1} : [e_i, e_1]=e_{i+1}, \quad 2\leq i \leq n-1.$$

Moreover, $[\mathfrak{s}_{n,2},\mathfrak{s}_{n,2}] = \mathfrak{n}_{n,1}$ and $\operatorname{Ann}_{[\mathfrak{s}_{n,2},\mathfrak{s}_{n,2}]}([\mathfrak{s}_{n,2},\mathfrak{s}_{n,2}]) = \langle e_n\rangle.$ 
In the following proposition we give the description of $\frac 1 2$-derivations of the algebra $\mathfrak{s}_{n,2}.$

\begin{proposition}  
Any $\frac 1 2$-derivation $\varphi$ of the algebra $\mathfrak{s}_{n,2}$
has the form
$$\varphi(x_1) = \alpha x_1 + (n-2)\beta e_n, \quad 
\varphi(x_2) = \alpha x_2 + \beta e_n, \quad 
\varphi(e_k) = \alpha e_k, \quad 1 \leq k  \leq n.$$
\end{proposition}

\begin{thrm} Let $(\mathfrak{s}_{n,2}, \cdot, [-,-])$ be a transposed Poisson algebra structure defined on the Lie
algebra $\mathfrak{s}_{n,2}.$ 
Then, up to isomorphism, there is only one non-trivial transposed Poisson algebra structure on $\mathfrak{s}_{n,2}$. It is given by
\begin{equation*}
 x_1 \cdot x_1 = (n-2)^2 e_n, \quad x_2 \cdot x_1 =  (n-2) e_n, \quad x_2 \cdot x_2 = e_n.   
\end{equation*}
This structure is non-Poisson.

\end{thrm}

\begin{proof} 
Let $(\mathfrak{s}_{n,2}, \cdot, [-,-])$ be a transposed Poisson algebra structure defined on the Lie
algebra $\mathfrak{s}_{n,2}.$ Then for any element of 
$x\in \mathfrak{s}_{n,2},$ the operator of multiplication
$\varphi_x(y) = x \cdot y$ is a  $\frac 1 2$-derivation
of $(\mathfrak{s}_{n,2}, [-,-])$. 
Hence,  we have that

{\small  

\begin{longtable}{lcl lcl lcl} 
$\varphi_{x_1}(x_1)$&$ = $&$\alpha_{n+1}x_1 + (n-2)\beta_{n+1}e_n,$ & 
$\varphi_{x_1}(x_2)$&$ = $&$ \alpha_{n+1} x_2 + \beta_{n+1} e_n,$ & 
$\varphi_{x_1}(e_k)$&$ =$&$ \alpha_{n+1} e_k,$\\

$\varphi_{x_2}(x_1) $&$=$&$ \alpha_{n+2}x_1 + (n-2)\beta_{n+2}e_n,$ &  
$\varphi_{x_2}(x_2) $&$=$&$ \alpha_{n+2} x_2 + \beta_{n+2} e_n,$ & 
$\varphi_{x_2}(e_k) $&$=$&$ \alpha_{n+2} e_k,$\\

$\varphi_{e_i}(x_1) $&$=$&$ \alpha_{i} x_1 + (n-2)\beta_ie_n,$ & 
$\varphi_{e_i}(x_2) $&$=$&$ \alpha_i x_2 + \beta_i e_n,$ & 
$\varphi_{e_i}(e_k) $&$=$&$ \alpha_i e_k.$
\end{longtable}}

Then from $\alpha_i e_j = \varphi_{e_i}(e_j) = e_i \cdot  e_j = e_j \cdot  e_i  =  \varphi_{e_j}(e_i) = \alpha_j e_i,$ we have that $\alpha_i =0$ for $1 \leq i \leq n.$
The last  implies $e_i \cdot e_j =0 $ for all $1 \leq i, j \leq n.$
Next from 
\begin{longtable}{llllllllllllll}
    $\alpha_{n+1} e_1 $&$=$&$ \varphi_{x_1}(e_1) $&$=$&$ e_1 \cdot  x_1$&$ =$&$ x_1 \cdot  e_1 $&$ =$&$ \varphi_{e_1}(x_1) $&$= $&$(n-2)\beta_1 e_n,$\\
 $\alpha_{n+2} e_1 $&$=$&$ \varphi_{x_2}(e_1)$&$=$&$ e_1 \cdot  x_2$&$ =$&$ x_2 \cdot  e_1 $&$ =$&$ \varphi_{e_1}(x_2) $&$=$&$ \beta_1 e_n,$
 \end{longtable}
we have that $\alpha_{n+1} =\alpha_{n+2} =0.$
After that, from 
\begin{longtable}{lllllllllllll}
    $0$&$=$&$\alpha_{n+2} e_i $&$=$&$ \varphi_{x_2}(e_i)$&$ = $&$e_i \cdot  x_2 $&$=$&$ x_2 \cdot  e_i  $&$=$&$ \varphi_{e_i}(x_2) $&$=$&$ \beta_i e_n,$    
\end{longtable}we get that $\beta_i =0$ for $1 \leq i \leq n$
and $e_i \cdot x_1=e_i\cdot x_2=0.$
Finally, considering 
\begin{longtable}{llllllllllllllllllllll}
    $(n-2)\beta_{n+2}e_n $&$=$&$ \varphi_{x_2}(x_1) $&$=$&$ x_2 \cdot  x_1 $&$=$&$ x_1 \cdot x_2  $&$=$&$ \varphi_{x_1}(x_2) $&$=$&$ \beta_{n+1} e_n,$
\end{longtable}
we get that $\beta_{n+1} =(n-2)\beta_{n+2}.$ Thus, we obtain
\begin{longtable}{lllllllllll}
$x_1 \cdot x_1 $&$= $&$(n-2)^2 \gamma e_n,$ & 
$x_2 \cdot x_1 $&$= $&$ (n-2) \gamma e_n,$ & 
$x_2 \cdot x_2 $&$=$&$ \gamma e_n.$
\end{longtable}

Since we are interested only in non-trivial transposed Poisson algebra structures, we can suppose $\gamma \neq 0$ and via automorphism 
\begin{longtable}{llllllllll}
$\phi(x_1) $&$= $&$x_1,$ &$\phi(x_2) = x_2,$ & $\phi(e_1) = e_1,$ &  
$\phi(e_i) = \gamma^{-1}{e_i},$ \ & $2 \leq i \leq n,$
\end{longtable}
we conclude that these transposed Poisson algebra structures are isomorphic to the case  $\gamma =1.$
\end{proof}

\subsection{Transposed Poisson structures on  solvable Lie algebras with Heisenberg nilpotent radical}
All solvable Lie algebras with Heisenberg nilpotent radical were found in  \cite{rw93}.
In this subsection, we consider $(3n+2)$-dimensional solvable Lie algebra $L_{n,n+1}$ with the following multiplication table: 
\begin{longtable}{lcllcllcll}
$[e_{n+i}, e_i]$&$=$&$e_{2n+1},$ &    $[e_i, x_{i}] $&$ =$&$ e_i,$  & $[e_{n+i}, x_{i}] $&$ =$&$ -e_{n+i},$  & $1 \leq i \leq n,$\\
& & & $[e_i, x_{n+1}] $&$=$&$ e_i,$  & $[e_{2n+1}, x_{n+1}] $&$= $&$e_{2n+1},$ & $1 \leq i \leq n.$
\end{longtable}

Note that the nilpotent radical of this algebra is the $(2n+1)$-dimensional 
Heisenberg algebra ${\rm H}_n$ with a basis $\{e_1, e_2, \dots, e_{2n+1}\}.$ 
Here $[L_{n,n+1},L_{n,n+1}] = {\rm H}_n$ and $\operatorname{Ann}_{{\rm H}_n}({\rm H}_n) = \langle e_{2n+1}\rangle.$
In the following proposition we give the description of $\frac 1 2$-derivations of the algebra $L_{n,n+1}.$

\begin{proposition} \label{prop-Heisenberg} 
Any $\frac 1 2$-derivation $\varphi$ of the algebra $L_{n,n+1}$
has the form
{\small \begin{longtable}{lcll lcll lcl}
$\varphi(e_k)$&$ =$&$ \alpha e_k, $&$ 1 \leq k  \leq 2n+1; $&$ 
\varphi(x_k) $&$= $&$\alpha x_k, $&$ 1 \leq k  \leq n; $&$ 
\varphi(x_{n+1}) $&$= $&$\alpha x_{n+1} + \beta e_{2n+1}.$
\end{longtable}}
\end{proposition}

\begin{thrm} 
Let $(L_{n,n+1}, \cdot, [-,-])$ be a transposed Poisson algebra structure defined on the Lie
algebra $L_{n,n+1}.$ 
Then, up to isomorphism, there is only one non-trivial transposed Poisson algebra structure on $L_{n,n+1}$. It is given by
 
$$x_{n+1} \cdot x_{n+1} =  e_{2n+1}.$$

This structure is non-Poisson.
\end{thrm}

\begin{proof} 
Let $(L_{n,n+1}, \cdot, [-,-])$ be a transposed Poisson algebra structure defined on the Lie
algebra $L_{n,n+1}.$ 
Then for any element of $x\in L_{n,n+1},$ 
we have the multiplication operator  
$\varphi_x(y) = x \cdot y$ is a  $\frac 1 2$-derivation of     $(L_{n,n+1}, [-,-]).$ Hence, we have that
{\small \begin{longtable}{lcll lcl}
$\varphi_{e_i}(e_k)$&$ = $&$\alpha_i e_k,$ & $1 \leq k  \leq 2n+1;$  \\
$\varphi_{e_i}(x_k)$&$ = $&$\alpha_i x_k,$ & $1 \leq k  \leq n+1;$ & $\varphi_{e_i}(x_{n+1}) $&$=$&$ \alpha_{i} x_{n+1} + \beta_ie_{2n+1},$ \\
$\varphi_{x_j}(e_k)$&$ =$&$ \alpha_{2n+1+j}e_k,$ & $1 \leq k  \leq 2n+1;$\\
$\varphi_{x_j}(x_k) $&$=$&$ \alpha_{2n+1+j} x_k,$ & $1 \leq k  \leq n;$ & 
$\varphi_{x_j}(x_{n+1}) $&$= $&$\alpha_{2n+1+j}x_{n+1} + \beta_{2n+1+j}e_{2n+1}.$
\end{longtable}}

Then from 
\begin{longtable}{lclclclclclc}
    $\alpha_i e_k $&$= $&$\varphi_{e_i}(e_k)$&$ =$&$ e_i \cdot  e_k $&$=$&$ e_k \cdot  e_i  $&$= $&$ \varphi_{e_k}(e_i) $&$=$&$ \alpha_k e_i,$    
\end{longtable}we have that $\alpha_i =0$ for $1 \leq i \leq 2n+1.$ Next from 
\begin{longtable}{lclclclclclc}
$\alpha_{2n+1+j} e_1 $&$=$&$ \varphi_{x_j}(e_1)$&$ =$&$ e_1 \cdot  x_j $&$=$&$ x_j \cdot  e_1  $&$=$&$ \varphi_{e_1}(x_j) $&$=$&$ 0,$
\end{longtable}
we have that $\alpha_{2n+1+j} =0$ for $1 \leq j \leq n.$ 
After then considering 
\begin{longtable}{lclclclclcl}
$\alpha_{3n+2} e_i$&$ =$&$ \varphi_{x_{n+1}}(e_i)$&$ =$&$ e_i \cdot  x_{n+1} $&$=$&$ x_{n+1} \cdot  e_i  $&$=$&$ \varphi_{e_i}(x_{n+1})$&$ =$&$ \beta_i e_{2n+1},$
\end{longtable}
we derive that $\alpha_{3n+2}=0$ and $\beta_i =0$ for $1 \leq i \leq 2n+1.$
Finally, considering 
\begin{longtable}{lclclclclcl}
$\beta_{2n+1+j}e_{2n+1} $&$=$&$ \varphi_{x_j}(x_{n+1}) $&$=$&$ x_j \cdot  x_{n+1} $&$ =$&$ x_{n+1} \cdot x_j  $&$= $&$\varphi_{x_{n+1}}(x_j) $&$= $&$0,$
\end{longtable} we get that $\beta_{2n+1+j} = 0$ for $1 \leq j \leq n$ and $x_i \cdot x_j = 0,$ $ 1 \leq i\leq n+1,$ $1 \leq j\leq n.$
Thus, we have that
$$ \quad x_{n+1} \cdot x_{n+1} = \gamma e_{2n+1}.$$

Since we are interested only in non-trivial transposed Poisson algebra structures, we have that $\gamma \neq 0$ and via  automorphism 
{\small \begin{longtable}{lclclclclclc}
$ \phi(e_i)$&$ =$&$ e_i,$ \ & $1 \leq i \leq n;$
     & $\phi(e_j) $&$= $&$\gamma^{-1}{e_j} ,$ \ & $n+1 \leq j \leq 2n+1;$ &
     $\phi(x_k) = x_k,$ \ & $1 \leq k \leq n+1,$
\end{longtable}}
we can suppose that $\gamma =1.$
\end{proof}

\subsection{Transposed Poisson structures on 
solvable Lie algebras with abelian nilpotent radical}

Now we consider solvable Lie algebras with abelian nilpotent radical and maximal complementary vector space
(these algebras were found in \cite{nw94}).
It is known that the maximal dimension of complementary space for solvable Lie algebras with $n$-dimensional abelian nilpotent radical is equal to $n.$ Moreover, up to isomorphism there exists only one such solvable Lie algebra with the following multiplications:
$$L_n: \left[e_i, x_i\right]=e_i, \quad 1 \leq i \leq n,$$
where $\{e_1,   \dots, e_n, x_1,  \dots, x_n\}$ is a basis of $L_n.$
In the following proposition we give the description of $\frac 1 2$-derivations of the algebra $L_{n}.$
\begin{proposition} Any $\frac 1 2$-derivation $\varphi$ of the algebra $L_{n}$
has the form
\begin{longtable}{lllllllll}
$\varphi(e_i) $&$=$&$ \alpha_i e_i,$  \  
& $\varphi(x_i) $&$= $&$ \alpha_i x_i + \beta_{i} e_i,$ \ $ 1 \leq i  \leq n.$
\end{longtable}
\end{proposition}

\begin{thrm}\label{thmL_n} Let $(L_n, \cdot, [-,-])$ be a transposed Poisson algebra structure defined on the Lie
algebra $L_{n}.$ Then the multiplication  of $(L_n, \cdot)$ has the following form:
$$e_i\cdot x_i=\mu_{i}e_{i}, \quad  x_i\cdot x_i = \mu_{i}x_i + \tau_{i}e_{i}, \quad  1 \leq i \leq n.$$
\end{thrm}

\begin{proof} $(L_n, \cdot, [-,-])$ be a transposed Poisson algebra structure defined on the Lie
algebra $L_{n}.$ Then for any element of $x\in L_n,$ we have that operator of multiplication
$\varphi_x(y) = x \cdot y$ is a  $\frac 1 2$-derivation. Hence,
for $ 1 \leq i,k  \leq n$ we derive
\begin{longtable}{lcll lcll}
$\varphi_{e_i}(e_k) $&$=$&$ \alpha_{i,k} e_k,$ &
$ \varphi_{e_i}(x_k) $&$= $&$\alpha_{i,k} x_k + \beta_{i,k}e_{k},$\\
$\varphi_{x_i}(e_k) $&$=$&$ \gamma_{i,k} e_k,$&
$ \varphi_{x_i}(x_k) $&$=$&$ \gamma_{i,k}x_k + \delta_{i,k}e_{k}.$
\end{longtable}


Thus, we have that
$$e_i\cdot e_k = 0, \quad 1 \leq i,k  \leq n; 
\quad e_j\cdot x_k=x_j\cdot x_k =0, \quad  1 \leq j\neq k \leq n.$$ 

Therefore, non-zero multiplications of the algebra $(L_n, \cdot)$ have the  form:
$$e_i\cdot x_i=\beta_{i,i}e_{i}, \quad x_i\cdot x_i = \beta_{i,i}x_i + \delta_{i,i}e_{i}, \quad  1 \leq i  \leq n.$$
\end{proof}

 \begin{remark} From Theorem \ref{thmL_n} we obtain that transposed Poisson algebra structure defined on the Lie
algebra $L_{n}$ is a direct sum of two-dimensional transposed Poisson algebras ${\rm TP}_i^{\mu_i, \tau_i},$ where
$${\rm TP}_i^{\mu_i, \tau_i}: \quad  
\left[e_i, x_i\right]=e_i, \quad e_i\cdot x_i=\mu_i e, \quad  x_i\cdot x_i = \mu_i x_i  + \tau_i e_i.$$
It is not difficult to see that any transposed Poisson algebra from the class ${\rm TP}_i^{\mu_i, \tau_i}$ is isomorphic one of the following non-isomorphic algebras
\begin{center}
    ${\rm TP}_1^\mu = {\rm TP}_i^{\mu, 0}$ 
    or ${\rm TP}_2={\rm TP}_i^{0, 1}.$
\end{center}

Therefore, we obtain that transposed Poisson algebra structure defined on the Lie algebra $L_{n}$ is a direct sum of some copies of algebras 
${\rm TP}_1^{\mu}$ and ${\rm TP}_2.$ 
 \end{remark}

\section{Transposed Poisson structure on perfect Lie algebras}

In this section, we consider transposed Poisson structure for some finite-dimensional Lie algebras which is a semi-direct sum of semisimple and solvable algebras, i.e., $\mathfrak{L} = \mathfrak{s} \ltimes \mathfrak{r}$. More precisely, we consider Lie algebras with three-dimensional simple parts and the solvable radical is an irreducible representation.  
It is known that if the  simple algebra $\mathfrak{sl}_2$ has an $(m+1)$-dimensional irreducible representation  $\mathfrak{r},$ where $m \geq 2$ (see, \cite{jac62}), 
then we have the algebra $\mathfrak L^m = \mathfrak{sl}_2 \ltimes \mathfrak r$ with  the following multiplication: 
{\small 
\begin{longtable}{lcllcllcl} 
$[e,f]$&$=$&$h,$ & $[h,e]$&$=$&$2e,$ & $[f,h]$&$=$&$2f,$\\  
$[x_k,h]$&$=$&$(2k-m)x_k,$ & \multicolumn{6}{l}{$0 \leq k \leq m ,$}\\
$[x_k,f]$&$=$&$x_{k+1},$  & \multicolumn{6}{l}{$0 \leq k \leq m-1,$} \\
 $[x_k,e]$&$=$&$k(m+1-k)x_{k-1},$ & \multicolumn{6}{l}{$1 \leq k \leq m.$}
 \end{longtable}}

\begin{proposition} \label{prop29}  If  $m> 2$, then any $\frac 1 2$-derivation of $\mathfrak L^m$ is trivial.
 \end{proposition}

\begin{proof} Let $\varphi$ be a $\frac 1 2$-derivation of $\mathfrak L^m.$ Put
\begin{longtable}{llllll}
$\varphi(e)$&$ =$&$ \alpha^1_1 e + \alpha^1_2 f + \alpha^1_3 h + \sum\limits_{k=0}^m\beta^1_{k}x_k,$ & 
$\varphi(f) $&$=$&$ \alpha^2_1 e + \alpha^2_2 f + \alpha^2_3 h + \sum\limits_{k=0}^m\beta^2_{k}x_k.$
\end{longtable}

Then from $\varphi(h) = \varphi([e,f]) = \frac 1 2 \big([\varphi(e),f] + [e, \varphi(f)]\big),$ we have that
{\small \begin{longtable}{lll}
    $\varphi(h)$&$ = $&$-\alpha^2_3 e - \alpha^1_3 f + \frac{\alpha^1_1+\alpha^2_2} 2 h - \frac m 2 \beta^2_{1}x_0  + \frac 1 2 \sum\limits_{k=1}^{m-1}\Big(\beta^1_{k-1} - (k+1)(m-k)\beta^2_{k+1} \Big) x_{k} + \frac 1 2 \beta^1_{m-1}x_{m}.$
    \end{longtable}}

Consider
{\small
\begin{longtable}{lcl}
$\varphi([h,e])$&$ = $&$\frac 1 2 \big([\varphi(h), e] + [h, \varphi(e)]\big)=$\\
&$=$& 
$\frac 1 2 \Big([-\alpha^2_3 e - \alpha^1_3 f + \frac{\alpha^1_1+\alpha^2_2} 2 h - \frac m 2 \beta^2_{1}x_0  +  \sum\limits_{k=1}^{m-1}\frac{\beta^1_{k-1} - (k+1)(m-k)\beta^2_{k+1}}{2} x_{k} + \frac{ \beta^1_{m-1}}{2}x_{m}, e] +$\\

\multicolumn{3}{r}{$ +[h, \alpha^1_1 e + \alpha^1_2 f + \alpha^1_3 h + \sum\limits_{k=0}^m\beta^1_{k}x_k]\Big)$}\\

&$=$&$\frac{\alpha^1_3} 2 h + \frac{\alpha^1_1+\alpha^2_2} 2 e + \sum\limits_{k=1}^{m-1}\frac{(\beta^1_{k-1} - (k+1)(m-k)\beta^2_{k+1})k(m+1-k)}{4} x_{k-1}+$\\

\multicolumn{3}{r}{$+\frac{m \beta^1_{m-1}}{4} x_{m-1} + \alpha^1_1 e - \alpha^1_2 f +    \sum\limits_{k=0}^m\frac{\beta^1_{k}(m-2k)}{2} x_k$}\\

&$=$&$\frac{3\alpha^1_1+\alpha^2_2} 2 e - \alpha^1_2 f + \frac{\alpha^1_3} 2 h +  \sum\limits_{k=0}^{m-2}\frac{(\beta^1_{k}- (k+2)(m-k-1)\beta^2_{k+2} )(k+1)(m-k)}{4} x_{k}+$\\
\multicolumn{3}{r}{$+\frac {m \beta^1_{m-1}}{4}x_{m-1} +  \sum\limits_{k=0}^m\frac{\beta^1_{k}(m-2k)}{2} x_k.$}
\end{longtable}
}

On the other hand, 
\begin{longtable}{lcllcllcl}
    $\varphi([h,e]) $&$ = $&$2 \varphi(e) $&$=$&$  2\alpha^1_1 e + 2\alpha^1_2 f + 2\alpha^1_3 h + 2\sum\limits_{k=0}^m\beta^1_{k}x_k.$
\end{longtable}
Comparing coefficients at the basis elements we have
$$\alpha^1_1= \alpha^2_2, \quad \alpha^1_2=0, \quad \alpha^1_3=0, \quad \beta_{1, m}=0, \quad \beta^1_{m-1}=0,$$
and (for  $0 \leq k \leq m-2$):
\begin{equation}\label{eq8}\big((k+1)(m-k) +2m -4k -8\big) \beta_{1,k}= (k+2)(k+1)(m-k)(m-k-1)\beta^2_{k+2}.\end{equation}

Let us now consider 
{\small \begin{longtable}{lcl}
$\varphi([f,h]) $&$= $&$
\frac 1 2 \big([\varphi(f), h] + [f, \varphi(h)]\big)$\\
&$=$&$\frac 1 2\big[\alpha^2_1 e + \alpha^1_1 f + \alpha^2_3 h +\sum\limits_{k=0}^m\beta_{2,k}x_k, h\big] +$\\

\multicolumn{3}{r}{$+\frac 1 2 \big[f, -\alpha^2_3 e + \alpha^1_1 h - \frac m 2 \beta^2_{1}x_0  +   
\sum\limits_{k=1}^{m-1}\frac{\beta^1_{k-1} - (k+1)(m-k)\beta^2_{k+1} }{2} x_{k} + \frac{\beta^1_{m-1}}{2}x_{m}\big]$}\\

&$=$&$-\alpha^2_1 e + \alpha^1_1 f +   
\sum\limits_{k=0}^m\frac{(2k-m)\beta_{2,k}}{2}x_k+ \frac{\alpha^2_3} 2 h + \alpha^1_1 f + \frac{m \beta^2_{1}}{4}x_1  -   \sum\limits_{k=2}^{m}\frac{\beta_{1,k-2} - k(m-k+1)\beta^2_{k}}{4} x_{k}.$
\end{longtable}}

On the other hand,
\begin{longtable}{lclclclclc}
$\varphi([f,h])$&$ =$&$ 2 \varphi(f)$&$ = $&$ 2\alpha^2_1 e + 2\alpha^1_1 f + 2\alpha^2_3 h + 2\sum\limits_{k=0}^m\beta^2_{k}x_k.$
\end{longtable}

Comparing coefficients at the basis elements we obtain
$$\alpha^2_1=0, \quad \alpha^2_3=0, \quad \beta_{2, 0}=0, \quad \beta_{2, 1}=0,$$ 
and
\begin{equation}\label{eq9}\beta_{1, k-2} = \big((k-2)m - k^2+5k-8\big)\beta_{2, k}, \quad 2 \leq k \leq m.\end{equation}

Then from \eqref{eq8} and \eqref{eq9} we have that $(m-2)(m+4)\beta_{2, k} =0,$ for $2 \leq k \leq m.$
Since $m\neq 2,$ then we have that $\beta_{2, k} =0,$ which implies $\beta_{1, k-2} =0$ for any  $2 \leq k \leq m.$

Thus, we have 
\begin{longtable}{lcllcllcl}
$\varphi(e) $&$=$&$\alpha^1_1 e,$&
$\varphi(f) $&$=$&$ \alpha^1_1 f,$&
$\varphi(h) $&$=$&$\alpha^1_1 h.$
\end{longtable}

Now put
\begin{longtable}{lcl}
$\varphi(x_0) $&$=$&$ \alpha^4_1 e + \alpha^4_2 f + \alpha^4_3 h + \sum\limits_{k=0}^m\beta^4_k x_k.$
\end{longtable}

From 
\begin{longtable}{lcl}
$0$&$=$&$2\varphi([x_0, e]) = [\varphi(x_0), e] + [x_0, \varphi(e)]  $\\
&$=$&$ [\alpha^4_1 e + \alpha^4_2 f + \alpha^4_3 h + \sum\limits_{k=0}^m\beta^4_k x_k, e] + [x_0, \alpha^1_1 e ]$\\
&$=$&$-\alpha^4_2 h + 2\alpha^4_3 e +
 \sum\limits_{k=1}^m k(m+1-k) \beta^4_k x_{k-1},$
\end{longtable}
we have that $\alpha^4_2 = \alpha^4_3  = \beta^4_k = 0,$ $1 \leq k \leq m.$
Considering the following relation 
\begin{longtable}{lclclcl}
$2m\varphi(x_0) $&$ =$&$ 2\varphi([h, x_0])  =  [\varphi(h), x_0] + [h, \varphi(x_0)] $\\
& $=$ &$[\alpha^1_1 h, x_0] + [h, \alpha^4_1 e + \beta^4_0 x_0]= 2\alpha^4_1 e + m(\alpha^1_1 + \beta^4_0) x_1,$
\end{longtable}
we have that $\alpha^4_1 = 0$ and $\beta^4_0 = \alpha^1_1.$ Thus, $\varphi(x_0) = \alpha^1_1x_0.$ 
Then from $\varphi(x_{k}) = \varphi([x_{k-1}, f]),$ we obtain that
$\varphi(x_k) = \alpha^1_1x_k$ for $1\leq k\leq m.$ 
Hence, any $\frac 1 2$-derivation of $\mathfrak L^m$ is trivial.
\end{proof}

Surprisingly, in the case of $m=2$, we obtain a different result.
\begin{proposition} \label{prop30}
Any $\frac 1 2$-derivation $\varphi$ of $\mathfrak L^2$  
has the following form
\begin{longtable}{lcllcllcl}
$\varphi(e) $&$= $&$\alpha e -2\beta x_0,$ & 
$\varphi(f) $&$= $&$ \alpha f + \beta x_2,$ & 
$\varphi(h) $&$= $&$ \alpha h -2 \beta x_1,$ \\
$\varphi(x_0) $&$= $&$\alpha x_0,$ & $\varphi(x_1) $&$ = $&$ \alpha x_1,$ & $\varphi(x_2) $&$= $&$\alpha x_2.$
\end{longtable}
 \end{proposition}

From the description of the $\frac 1 2$-derivation of the algebra $\mathfrak{L}^m$ we have the following corollary. 
\begin{corollary} There are no non-trivial transposed Poisson algebra structures defined on the algebra $\mathfrak L^m.$  \end{corollary}
\begin{proof} Since in the case of $m\neq 2,$  the algebra $\mathfrak L^m$ has only trivial $\frac 1 2$-derivation,  we obtain that there are no non-trivial transposed Poisson algebra structures for $m\neq 2.$ 

Thus, it is sufficient to prove the Corollary for $m=2.$
Then we have the multiplication of $\mathfrak L^2$ has the form 
\begin{longtable}{lcllcllcl} 
$[e,f] $&$=$&$h,$ & $[h,e]$&$=$&$2e,$ & $[f,h]$&$=$&$2f,$ \\
$[x_0,h]$&$=$&$-2x_0,$ & $[x_0,f]$&$=$&$x_1,$ & $[x_1,e]$&$=$&$2x_0,$ \\
$[x_2,h]$&$=$&$2x_2,$ & $[x_1,f]$&$=$&$x_2,$ & $ [x_2,e]$&$=$&$2x_1.$ \\
\end{longtable}

By Lemma \ref{glavlem},  we have that for any element   
$\ell \in \{e, f, h, x_0, x_1, x_2\}$, there is a related 
$\frac{1}{2}$-derivation $\varphi_{\ell}$ of $\mathfrak L^2$.
Thus, by Proposition \ref{prop30}, we derive that
$$\begin{array}{lll} \varphi_{\ell}(e) = \alpha_{\ell} e -2\beta_{\ell} x_0, & \varphi_{\ell}(f) = \alpha_{\ell} f + \beta_{\ell} x_2, & \varphi_{\ell}(h) = \alpha_{\ell} h -2 \beta_{\ell} x_1, \\[1mm]
\varphi_{\ell}(x_0) = \alpha_{\ell} x_0, & \varphi_{\ell}(x_1) = \alpha_{\ell} x_1, & \varphi_{\ell}(x_2) = \alpha_{\ell} x_2.\end{array}$$
 
Then from 
\begin{longtable}{lcllcllcllcl}
$\alpha_{x_i} x_j $&$= $&$\varphi_{x_i}(x_j)$&$ =$&$ x_i \cdot x_j $&$= $&$\varphi_{x_j}(x_i) = \alpha_{x_j} x_i,$
\end{longtable}
we have that $\alpha_{x_0}=\alpha_{x_1} = \alpha_{x_2} =0.$ 

Now, considering 
\begin{longtable}{lclclclclclclcl}
$\alpha_{e} x_1 $&$=$&$ \varphi_{e}(x_1) $&$=$ &$e \cdot x_1 $&$= $&$\varphi_{x_1}(e) $&$=$&$ -2\beta_{x_1} x_0,$\\ 
$\alpha_{f} x_0 $&$=$&$ \varphi_{f}(x_0) $&$=$&$ f \cdot x_0 $&$=$&$ \varphi_{x_0}(f) $&$=$&$ \beta_{x_0} x_2,$\\ 
$\alpha_{h} x_2 $&$=$&$ \varphi_{h}(x_2) $&$=$&$ h \cdot x_2 $&$=$&$ \varphi_{x_2}(h) $&$=$&$ -2\beta_{x_2} x_1,$
\end{longtable}
we obtain that $\alpha_{e}=\alpha_{f} = \alpha_{h} = \beta_{x_0} = \beta_{x_1} = \beta_{x_2}=0.$

Finally, from
\begin{longtable}{lclclclclclclcl}
$\beta_{e} x_2$ &$=$& $\varphi_{e}(f)$ &$=$ &$e \cdot f$ &$=$ &$\varphi_{f}(e) $&$=$&$ -2\beta_{f} x_0,$\\
$-2\beta_{e} x_1 $&$=$&$ \varphi_{e}(h) $&$= $&$e \cdot h $&$= $&$\varphi_{h}(e) $&$=$&$ -2\beta_{h} x_0,$
\end{longtable}
we get that $\beta_{e} = \beta_{f} = \beta_{h}=0.$ Thus, we conclude that there are no non-trivial transposed Poisson algebra structures on $\mathfrak L^2$.
\end{proof}

\begin{remark} It should be noted that the algebra $\mathfrak L^2$ is an algebra that admits non-trivial $\frac{1}{2}$-derivations, but
does not admit non-trivial transposed Poisson algebra structures.
Moreover, the algebra $\mathfrak L^2$ is a perfect algebra 
(i.e., $[\mathfrak L^2,\mathfrak L^2] =\mathfrak L^2$) 
and  admits non-trivial $\frac{1}{2}$-derivations.
The existence of algebras with this property gives a particular answer to \cite[Questions 2 and 5]{bfk23}.
    
\end{remark}

\end{document}